# Bounding causal effects with an unknown mixture of informative and non-informative missingness

### v3

@@ 5.2 Estimation
The remaining challenge is estimating $\ell_{2}$ and $u_{2}$. These functionals are not pathwise-differentiable due to the presence of the indicator functions ($\mathbb{I}\{\mu_{1}(\mathbf{X})\geq\mu_{0}(\mathbf{X})\}$ and $\mathbb{I}\{\mu_{1}(\mathbf{X})<\mu_{0}(\mathbf{X})\}$) (branson2023causal) and therefore have no influence functions. One option is to replace the indicator functions with smooth approximations, such as the centered normal cumulative distribution function $\Phi_{\epsilon}$, where $\epsilon$ parameterizes the standard deviation. In other words, instead of targeting $u_{2}=\tilde{\Psi}-\mathbb{E}[\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\mathbb{I}\{\mu_{1}(\mathbf{X})\geq\mu_{0}(\mathbf{X})\}]$, we may instead target $u_{2,\epsilon}=\tilde{\Psi}-\mathbb{E}[\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\Phi_{\epsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))]$. While these estimands are not identical, the difference can be made arbitrarily small by decreasing the parameter $\epsilon$. [^6] However, as we decrease $\epsilon$, we may increase the variance of the corresponding estimator yang2018asymptotic; branson2023causal, leading to a bias-variance tradeoff. Proposition 11 gives the influence function for $u_{2,\epsilon}$.
@@
For fixed $\epsilon>0$, the uncentered influence function for $u_{2,\epsilon}$ is
$\varphi_{u_{2}}({\bf O};\eta)$ $=\left(\frac{\mathbb{I}\{C=0,A=1\}}{(1-\pi_{1}(\mathbf{X}))e_{1}(\mathbf{X})}-\frac{\mathbb{I}(C=0,A=0)}{(1-\pi_{0}(\mathbf{X}))e_{0}(\mathbf{X}))}\right)\{Y-\mu_{A}(\mathbf{X})\}\pi_{0}(\mathbf{X})\Phi_{\epsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))$
$+\frac{\mathbb{I}\{A=0\}}{e_{0}(\mathbf{X})}\{C-\pi_{0}(\mathbf{X})\}[\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})]\Phi_{\epsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))$
$+[\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})]\pi_{0}(\mathbf{X})\Phi_{\epsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))$
$+[\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})]\pi_{0}(\mathbf{X})\phi_{\epsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))[\varphi_{1,1}(\mathbf{X})-\varphi_{1,0}(\mathbf{X})-\mu_{1}(\mathbf{X})+\mu_{0}(\mathbf{X})].$ (5)
where $\phi_{\epsilon}$ is the probability density function corresponding to $\Phi_{\epsilon}$.
@@
The influence function for $\ell_{2,\epsilon}$ is similar but reverses the signs of $\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})$ in the functions $\Phi_{\epsilon}$ and $\phi_{\epsilon}$, as well as in the term $\varphi_{1,1}({\bf O})-\varphi_{1,0}({\bf O})-\mu_{1}(\mathbf{X})+\mu_{0}(\mathbf{X})$. The full expression is available in Appendix 9.
To obtain an estimator of $u_{2,\epsilon}$, we plug in estimates of the nuisance components and compute the empirical average of these quantities. Theorem 2 in Appendix 9 establishes the conditions under which this estimator is root-$n$ consistent and asymptotically normal.
As an alternative option, one can estimate $u_{2}$ ($\ell_{2}$) by eliminating the final line 5, and replacing $\Phi_{\epsilon}$ with $\mathbb{I}(\mu_{1}(\mathbf{X})\geq\mu_{0}(\mathbf{X}))$ ($\mathbb{I}(\mu_{1}(\mathbf{X})<\mu_{0}(\mathbf{X}))$). This estimator will also be root-n consistent and asymptotically normal under similar conditions to Theorem 2, which we detail in Theorem 3 in Appendix 9, and an additional so-called “margin condition” that essentially requires the density of $\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})$ to be bounded. This condition is likely reasonable for many applications, though would be violated if, for example, the sharp null held with respect to the true conditional treatment effects, and there were no differential dropout, so that $\mu_{1}(\mathbf{x})=\mu_{0}(\mathbf{x})$ for all $\mathbf{x}$.
@@ 6 Simulations
We assess the theoretical properties of our proposed estimators through a simulation study. We begin by generating a population of size $N=2,000,000$. For each individual i=1,,Ni=1,...,N, we draw a single covariate independently $X_{i}\sim U[-3,3]$ and generate $A_{i},C_{i},Y_{i}$ according to the nuisance functions depicted below in Figure 2. As illustrated, the propensity score model $e(x)$, the outcome risk $\mu_{a}(x)$, and the missingness probabilities $\pi_{a}(x)$ are all increasing functions of $x$. A complete description of the data-generating process is provided in Appendix 11.
@@
Next, we take 1,000 random samples of size $n=1,000$. Following the simulation framework of kennedy2023towards, for each nuisance parameter $\eta_{j}$, we simulate an estimated function $\hat{\eta}_{j}$ by introducing stochastic perturbations: $\hat{\eta}_{j}=\text{expit}(\text{logit}(\eta_{j})+\mathcal{N}(C_{1j}n^{-\alpha},C_{2j}n^{-2\alpha}))$ for constants $C_{1j}$ and $C_{2j}$ that are specific to each nuisance function and parameter to ensure that we generate estimates that satisfy $\|\hat{\eta}-\eta\|=\mathcal{O}_{p}(n^{-\alpha})$ in the $L_{2}$ norm. Using the simulated estimates $\hat{\eta}$, we construct the estimated influence functions and take the empirical average and variance to construct confidence intervals.
We consider two scenarios for the convergence rate parameter $\alpha$: $0.3$ and $0.1$. In the first case, the convergence rate satisfies the conditions of Proposition 6, and therefore the estimators are theoretically fast enough to achieve root-$n$ consistency and asymptotic normality. However, the second case – $\alpha=0.1$ – does not satisfy the rate conditions of Proposition 6, and therefore, we should not expect the estimator to perform well. If we believed that this was the convergence rate for an optimal non-parametric estimator, we would not be able to obtain a root-n convergent estimator without correctly specifying the nuisance component models.
@@
To evaluate the practical performance of our estimators, we conduct a second simulation study where we estimate the nuisance functions $\hat{\eta}_{j}$ directly using sample splitting and SuperLearner to stack a generalized linear model and random forest for samples of size $n=2,000$. Given the relatively simple data-generating processes, we expect the nuisance estimators to achieve convergence rates of at least $n^{-1/4}$, yielding similar results to the $\alpha=0.3$ setting above. The final column in Table 1 displays these results and confirms this expectation. The estimators demonstrate minimal bias and confidence intervals with coverage rates close to the nominal $95\%$ level.
@@ 7 Application
Antipsychotic drugs (APs) are a mainstay in the treatment of schizophrenia. However, some APs are associated with safety risks, including A1c disregulation, type 2 diabetes, and other cardiometabolic morbidity. Randomized studies comparing APs on safety events that like type 2 diabetes take months or years to develop, are virtually non-existent; similarly, there is very limited randomized evidence on the effectiveness of APs among elderly individuals with schizophrenia, a population who may plausibly respond differently to these drugs relative to younger patients. Consequently, clinicians must rely on evidence from observational studies to guide their clinical decision-making. Health insurance administrative databases, especially when paired with state-of-the-art causal inference methods, offer the promise of helping inform the relative effectiveness and safety in usual care populations. However, the potential for informative missingness is particularly acute in many of these databases because reasons for missingness are often not ascertained or communicated to the researcher.
We examined data on $n=11,618$ elderly adults (age greater than 66) insured by Medicare Advantage diagnosed with schizophrenia who were relatively new users of APs identified in Inovalon Insights LLC’s Payer Sourced Claims Data (2017-2020). [For more details on inclusion criteria in a similar population, see agniel2024revisiting]. We studied the causal effect of each of five common APs compared with taking any other AP in our data: aripiprazole ($n=975$), haloperidol ($n=1,024$), olanzapine ($n=1,798$), quetiapine ($n=3,702$), and risperidone ($n=3,036$). People were assigned to a particular AP based on their first qualifying fill in the database. A total of 1,083 patients took other APs. We measured two outcomes that may play an important role in clinicians’ prescribing decisionmaking: diagnosis of type 2 diabetes (which informs AP safety) and inpatient hospitalization for any psychiatric condition (which informs AP effectiveness), both measured at 12 months after the initial AP fill. Missingness was common ($20\%$) for both outcomes in all AP groups.
Estimation of all nuisance components was undertaken using random forests, and cross-fitting with two folds to separate evaluation of the influence function from nuisance function estimation. To account for additional randomness due to sample splitting, we repeated sample-splitting and estimation 11 times, using different seeds each time. We took estimates to be the median of the 11 estimates, and we computed standard errors using the approach in chernozhukov2018double, which accounts for additional variability due to different seeds across the 11 estimates.
@@
Antipsychotic $\widetilde{\Phi}$} [CI], diabetes $\widetilde{\Phi}$} [CI], hospitalization
@@
This most general analysis allows informative missingness to either increase or decrease the risk of the outcomes. In this elderly population with schizophrenia, we are most concerned about missingness that is due to causes that might increase their risk of these adverse outcomes. We may encode this using Assumption 7.c, and we can use the bounds from Proposition 3 to provide more narrow bounds. Based on this proposition, we have a negative upper bound for the ATE ($u_{0}^{\text{pos}}<0$) as long as the amount of informative missingness in the treated group ($\delta_{1u}$) is not too high. For example, $u_{0}^{\text{pos}}<0$ for haloperidol’s effect on diabetes as long as no more than 40% of the missingness is informative. The estimates for aripiprazole’s and quetiapine’s effects on diabetes are slightly less robust, requiring no more than 21% and 17% informative missingness, respectively. The psychiatric inpatient hospitalization (effectiveness) analyses, on the other hand, are not robust to informative missingness at all. Even the largest magnitude effect on inpatient hospitalizations requires no more than 5% informative missingness.
Finally, we can examine what type of increased risk due to informative missingness would be required to explain away the naive estimate $\widetilde{\Psi}$}. Under Assumption 7.f, we require $\tau>-{\widetilde{\Psi}}/\mathbb{E}[\mu_{1}(\mathbf{X})\pi_{1}(\mathbf{X})]+1$, if $\Psi_{0}$ has any possibility of being non-negative. For the ATE of haloperidol on diabetes, we require informative missingness to increase diabetes risk by a factor of 10.7 to explain the naive ATE estimate. Similarly, aripiprazole requires $\tau>4.0$ – a still sizable increase – while quetiapine is less robust requiring a more modest increased risk of diabetes of a factor of $\tau>1.7$.
In short, unless the severity of informative missingness is quite high and assuming that there is no unmeasured confounding, we find the strongest evidence that haloperidol decreases the risk of diabetes among elderly adults with schizophrenia relative to all other drugs we study, a result that is consistent with results from poulos2023antipsychotics.
@@ 8 Discussion
We have proposed a novel framework for bounding treatment effects and conducting sensitivity analyses in the presence of a mixture of informative and non-informative missingness. Just as most observational studies likely suffer from some amount of unobserved confounding, most studies with missingness likely suffer from some amount of informative missingness. To address this, we have developed assumption-lean bounds that are broadly applicable in virtually any study with missingness and a binary outcome. We also provide a variety of additional assumptions that researchers may use to narrow these bounds or conduct sensitivity analyses in more general settings. Bounding approaches may be preferred when realistic values of sensitivity parameters can reasonably be specified, while “tipping point”-style sensitivity analyses may be preferred when researchers may only know broadly what values of sensitivity parameters are implausible.
However, a limitation of our proposed approach is that the estimation is not guaranteed to be constrained to the natural bounds of a bounded outcome. For example, in sensitivity analyses where one might wish to evaluate and reason about quantities like $\mathbb{E}[\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})]$, these quantities could be negative even if the outcome is binary. A straightforward adaptation of our current approach, which leverages a one-step estimator based on the influence function, would be to implement targeted minimum loss-based estimation (TMLE) (van2011targeted). TMLE could enforce the necessary constraints, ensuring that the estimators remain within plausible bounds.
We have also limited our work to the setting with a single time point at which missingness occurs. However, many practical studies face scenarios where missingness occurs at multiple time points, and outcomes may be observed at different intervals. Extending our approach to a longitudinal or survival context is conceptually straightforward but would introduce significant notational complexity and require additional sensitivity parameters. To make such tools practical, further assumptions linking informative missingness to changes in risk over time would be necessary. The development of a comprehensive system for handling these more complex settings is an important area for future research.
@@
Acknowledgments

### v4

@@ 5.2 Estimation
The remaining challenge is estimating $\ell_{2}$ and $u_{2}$. These functionals are not pathwise-differentiable due to the presence of the indicator functions ($\mathbb{I}\{\mu_{1}(\mathbf{X})\geq\mu_{0}(\mathbf{X})\}$ and $\mathbb{I}\{\mu_{1}(\mathbf{X})<\mu_{0}(\mathbf{X})\}$) Branson et al. (2023) and therefore have no influence functions. One option is to replace the indicator functions with smooth approximations, such as the centered normal cumulative distribution function $\Phi_{\varepsilon}$, where $\varepsilon$ parameterizes the standard deviation. In other words, instead of targeting $\ell_{2}=\tilde{\Psi}-\mathbb{E}[\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\mathbb{I}\{\mu_{1}(\mathbf{X})>\mu_{0}(\mathbf{X})\}]$, we may instead target $\ell_{2,\varepsilon}=\tilde{\Psi}-\mathbb{E}[\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\Phi_{\varepsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))]$. While these estimands are not identical, the difference can be made arbitrarily small by decreasing the parameter $\varepsilon$. [^6] However, as we decrease $\varepsilon$, we may increase the variance of the corresponding estimator Yang and Ding (2018); Branson et al. (2023), leading to a bias-variance tradeoff. Proposition 11 gives the influence function for $u_{2,\varepsilon}$.
@@
For a fixed $\varepsilon>0$, the uncentered influence function for $\ell_{2,\varepsilon}$ is
$\varphi_{\ell_{2}}({\bf O};\eta)$ $=\varphi_{1,1}({\bf O})-\varphi_{1,0}({\bf O})$
$-\left(\frac{\mathbb{I}\{C=0,A=1\}}{(1-\pi_{1}(\mathbf{X}))e_{1}(\mathbf{X})}-\frac{\mathbb{I}(C=0,A=0)}{(1-\pi_{0}(\mathbf{X}))e_{0}(\mathbf{X})}\right)\{Y-\mu_{A}(\mathbf{X})\}\pi_{0}(\mathbf{X})\Phi_{\varepsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))$ (6)
$-\frac{\mathbb{I}\{A=0\}}{e_{0}(\mathbf{X})}\{C-\pi_{0}(\mathbf{X})\}[\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})]\Phi_{\varepsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))$
$-[\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})]\pi_{0}(\mathbf{X})\Phi_{\varepsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))$
$-[\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})]\pi_{0}(\mathbf{X})\phi_{\varepsilon}(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\{\varphi_{1,1}({\bf O})-\varphi_{1,0}({\bf O})-\mu_{1}(\mathbf{X})+\mu_{0}(\mathbf{X})\}.$ (7)
where $\phi_{\varepsilon}$ is the probability density function corresponding to $\Phi_{\varepsilon}$.
@@
The influence function for $u_{2,\varepsilon}$ is similar but reverses the signs of $\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})$ in the functions $\Phi_{\varepsilon}$ and $\phi_{\varepsilon}$, as well as in the term $\varphi_{1,1}({\bf O})-\varphi_{1,0}({\bf O})-\mu_{1}(\mathbf{X})+\mu_{0}(\mathbf{X})$. The full expression is available in Appendix 9.
To obtain an estimator of $\ell_{2,\epsilon}$, we plug in estimates of the nuisance components and compute the empirical average of these quantities. Theorem 2 in Appendix 9 establishes the conditions under which this estimator is root-$n$ consistent and asymptotically normal.
As an alternative option, one can estimate $\ell_{2}$ ($u_{2}$) by eliminating the final line 7, and replacing $\Phi_{\epsilon}$ with $\mathbb{I}(\mu_{1}(\mathbf{X})>\mu_{0}(\mathbf{X}))$ ($\mathbb{I}(\mu_{1}(\mathbf{X})\leq\mu_{0}(\mathbf{X}))$). This estimator will also be root-n consistent and asymptotically normal under similar conditions to those given Theorem 2. We formalize this result in Theorem 3 in Appendix 9, invoking an additional “margin condition” that essentially requires the density of $\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})$ to be bounded. This condition is likely reasonable for many applications, though would be violated if, for example, the sharp null held with respect to the true conditional treatment effects, and there were no differential dropout, so that $\mu_{1}(\mathbf{x})=\mu_{0}(\mathbf{x})$ for all $\mathbf{x}$.
@@ 6 Simulations
We assess the theoretical properties of our proposed estimators through a simulation study. We begin by generating a population of size $N=2,000,000$. For each individual i=1,,Ni=1,...,N, we draw a single covariate independently $X_{i}\sim\mathcal{U}[-3,3]$ and generate $A_{i},C_{i},Y_{i}$ according to the nuisance functions depicted below in Figure 2. As illustrated, the propensity score model $e(x)$, the outcome risk $\mu_{a}(x)$, and the missingness probabilities $\pi_{a}(x)$ are all increasing functions of $x$. A complete description of the data-generating process is provided in Appendix 12.
@@
Next, we take 1,000 random samples of size $n=1,000$. Following the simulation framework of Kennedy (2023), for each nuisance parameter $\eta_{j}$, we simulate an estimated function $\hat{\eta}_{j}$ by introducing stochastic perturbations: $\hat{\eta}_{j}=\text{expit}(\text{logit}(\eta_{j})+\mathcal{N}(C_{1j}n^{-\alpha},C_{2j}n^{-2\alpha}))$ for constants $C_{1j}$ and $C_{2j}$ that are specific to each nuisance function and parameter to ensure that we generate estimates that satisfy $\|\hat{\eta}-\eta\|=\mathcal{O}_{p}(n^{-\alpha})$ in the $L_{2}(\mathbb{P})$ norm. Using the simulated estimates $\hat{\eta}$, we construct the estimated influence functions and take the empirical average and variance to construct confidence intervals.
We consider two scenarios for the convergence rate parameter $\alpha$: $0.3$ and $0.1$. In the first case, the convergence rate satisfies the conditions of Theorem 1, and therefore the estimators are theoretically fast enough to achieve root-$n$ consistency and asymptotic normality. However, the second case – $\alpha=0.1$ – does not satisfy the rate conditions of Theorem 1, and therefore, we should not expect the estimator to perform well. If we believed that this was the convergence rate for an optimal non-parametric estimator, we would not be able to obtain a root-n convergent estimator without correctly specifying the nuisance component models.
@@
To evaluate the practical performance of our estimators, we conduct a second simulation study where we estimate the nuisance functions $\hat{\eta}_{j}$ directly using sample splitting and SuperLearner to stack a generalized linear model and random forest for samples of size $n=2,000$. Given the relatively simple data-generating processes, we expect the nuisance estimators to achieve convergence rates greater than $n^{-1/4}$, yielding similar results to the $\alpha=0.3$ setting above. The final column in Table 1 displays these results and confirms this expectation. The estimators demonstrate minimal bias and confidence intervals with coverage rates close to the nominal $95\%$ level.
@@ 7 Application
Antipsychotic drugs (APs) are a mainstay in the treatment of schizophrenia. However, some APs are associated with safety risks, including A1c disregulation, type 2 diabetes, and other cardiometabolic morbidity. Randomized studies comparing APs on safety events that, like type 2 diabetes, take months or years to develop, are virtually non-existent; similarly, there is very limited randomized evidence on the effectiveness of APs among elderly individuals with schizophrenia, a population who may plausibly respond differently to these drugs relative to younger patients. Consequently, clinicians must rely on evidence from observational studies to guide their clinical decision-making. Health insurance administrative databases, especially when paired with state-of-the-art causal inference methods, offer the promise of helping inform the relative effectiveness and safety in usual care populations. However, the potential for informative missingness is particularly acute in many of these databases because reasons for missingness are often not ascertained or communicated to the researcher.
We examined data on $n=11,618$ elderly adults (age greater than 66) insured by Medicare Advantage diagnosed with schizophrenia who were relatively new users of APs identified in Inovalon Insights LLC’s Payer Sourced Claims Data (2017-2020). For more details on inclusion criteria in a similar population, see Agniel et al. (2024).
We studied the causal effect of each of five common APs compared with taking any other AP in our data: aripiprazole ($n=975$), haloperidol ($n=1,024$), olanzapine ($n=1,798$), quetiapine ($n=3,702$), and risperidone ($n=3,036$). People were assigned to a particular AP based on their first qualifying fill in the database. A total of 1,083 patients took other APs. We measured two outcomes that may play an important role in clinicians’ prescribing decision-making: diagnosis of type 2 diabetes (which informs AP safety) and inpatient hospitalization for any psychiatric condition (which informs AP effectiveness), both measured at 12 months after the initial AP fill. Missingness was common ($20\%$) for both outcomes in all AP groups.
Estimation of all nuisance components was undertaken using random forests, and cross-fitting with two folds to separate evaluation of the influence function from nuisance function estimation. To account for additional randomness due to sample splitting, we repeated sample-splitting and estimation 11 times, using different seeds each time. We took estimates to be the median of the 11 estimates, and we computed standard errors using the approach in Chernozhukov et al. (2018), which accounts for additional variability due to different seeds across the 11 estimates.
@@
Antipsychotic $\widetilde{\Psi}$} [CI], diabetes $\widetilde{\Psi}$} [CI], hospitalization
@@
This most general analysis allows informative missingness to either increase or decrease the risk of the outcomes. In this elderly population with schizophrenia, we are most concerned about missingness that is due to causes that might increase their risk of these adverse outcomes. We may encode this using Assumption 7.a, and we can use the bounds from Proposition 3 to provide more narrow bounds. Based on this proposition, we have a negative upper bound for the ATE ($u_{0}^{\text{pos}}<0$) as long as the amount of informative missingness in the treated group ($\delta_{1u}$) is not too high. For example, $u_{0}^{\text{pos}}<0$ for haloperidol’s effect on diabetes as long as no more than 40% of the missingness is informative. The estimates for aripiprazole’s and quetiapine’s effects on diabetes are slightly less robust, requiring no more than 21% and 17% informative missingness, respectively. The psychiatric inpatient hospitalization (effectiveness) analyses, on the other hand, are not robust to informative missingness at all. Even the largest magnitude effect on inpatient hospitalizations requires no more than 5% informative missingness.
Finally, we can examine what type of increased risk due to informative missingness would be required to explain away the naive estimate $\widetilde{\Psi}$}. Under Assumption 7.d, we require $\tau>-{\widetilde{\Psi}}/\mathbb{E}[\mu_{1}(\mathbf{X})\pi_{1}(\mathbf{X})]+1$, if $\Psi_{0}$ has any possibility of being non-negative. For the ATE of haloperidol on diabetes, we require informative missingness to increase diabetes risk by a factor of 10.7 to explain the naive ATE estimate. Similarly, aripiprazole requires $\tau>4.0$ – a still sizable increase – while quetiapine is less robust requiring a more modest increased risk of diabetes of a factor of $\tau>1.7$. [^7] These numbers are estimated; for simplicity, we do not quantify uncertainty around these estimates.
In short, unless the severity of informative missingness is quite high and assuming that there is no unmeasured confounding, we find the strongest evidence that haloperidol decreases the risk of diabetes among elderly adults with schizophrenia relative to all other drugs we study, a result that is consistent with findings from Poulos et al. (2023).
@@ 8 Discussion
We have proposed a novel framework for bounding treatment effects and conducting sensitivity analyses in the presence of a mixture of informative and non-informative missingness. Just as most observational studies likely suffer from some amount of unobserved confounding, most studies with missingness likely suffer from some amount of informative missingness. To address this, we have developed assumption-lean bounds that are broadly applicable in virtually any study with missingness and a bounded outcome. We also provide a variety of additional assumptions that researchers may use to narrow these bounds or conduct sensitivity analyses in more general settings. Bounding approaches may be preferred when realistic values of sensitivity parameters can reasonably be specified, while “tipping point”-style sensitivity analyses may be preferred when researchers may only know broadly what values of sensitivity parameters are implausible. In many applications, researchers may instead wish to estimate the bounds over a range of sensitivity parameters, which can both identify a tipping point (or tipping region) and provide a broader picture of how sensitive the bounds are to the assumed sensitivity parameter values.
A limitation of our proposed approach is that the estimation is not guaranteed to be constrained to the natural bounds of a bounded outcome. For example, in sensitivity analyses where one might wish to evaluate and reason about quantities like $\mathbb{E}[\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})]$, these quantities could be negative even if the outcome is binary. A straightforward adaptation of our current approach, which leverages a one-step estimator based on the influence function, would be to implement targeted minimum loss-based estimation (TMLE) Van der Laan et al. (2011). TMLE could enforce the necessary constraints, ensuring that the estimators remain within plausible bounds.
We have also limited our work to the setting with a single time point at which missingness occurs. However, many studies face scenarios where missingness occurs at multiple time points, and outcomes may be observed at different intervals. Extending our approach to a longitudinal or survival context is conceptually straightforward but would introduce significant notational complexity and require additional sensitivity parameters. To make such tools practical, further assumptions linking informative missingness to changes in risk over time would be necessary. The development of a comprehensive system for handling these more complex settings is an important area for future research.
@@
Acknowledgments: The authors are indebted to Ben Buzzee, Harvard Medical School, for creation of the antipsychotic dataset. The authors would also like to thank two anonymous reviewers and the Associate Editor for helpful comments, questions, and suggestions that substantially improved the quality of this manuscript. Finally, the authors used ChatGPT for assistance proofreading and Claude Code to write and document code. All mistakes are our own.